# 3D-Speaker-Toolkit: An Open-Source Toolkit for Multimodal Speaker Verification and Diarization

### v2

@@ I Introduction
The research of speaker representation learning has achieved remarkable progress in recent years [1, 2, 3, 4, 5]. Speaker representation learning captures the unique characteristics of a speaker in a compact form, and is extensively utilized across various tasks including speaker verification [3], speaker diarization [6], target speaker extraction [7], speaker-attributed automatic speech recognition [8], and other speech-related tasks [9]. Researchers have studied different training schemes to learn robust speaker representations, from fully-supervised leaning [3, 1] to self-supervised learning [12, 11, 10]. With the continuous emergence of open-source projects, speech-related technology has advanced in sophistication. Table I lists some well-known open-source speech toolkits. Initially, foundational speech processing toolkits such as Kaldi [13] serve as crucial resources for both researchers and industrial applications. More recently, ESPnet [14] and SpeechBrain [15] provide novice-friendly code designs and usage. Furthermore, VoxCeleb_Trainer [16], ASV-Subtools [17], and Wespeaker [18] offer open-source platforms for the speaker community to effortlessly build models.
@@ II-B1 Visual Module
[…]
@@
TABLE IV: Comparison between self-supervised learning models. “*” denotes the SSL frameworks in 3D-Speaker-Toolkit.
Model Embedding Extractor EER(%)
@@ III-B Multimodal Speaker diarization
Experiments based on the proposed multimodal diarization method are conducted on a self-collected video dataset, which includes 2 to 10 speakers. Strong baselines for audio-only diarization have been established using VBx [49] and Spectral clustering(SC) [50] methods. We use E2CP intergrated with spectral clustering method to leverage multimodal information to enhance speaker diarization. The comparison results are shown in Table V. Common speaker diarization metrics, including Diarization Error Rate(DER) [51], Jaccard Error Rate(JER) [52] and Concatenated Minimum-permutation Word Error Rate(CpWER) [53], are used. The results show that, compared to systems based solely on audio, incorporating visual or textual information both achieve notable performance improvements. Combining all three modalities of audio, visual, and textual data yields superior results on all metrics.

### v3

@@ I Introduction
The research of speaker representation learning has achieved remarkable progress in recent years [1, 2, 3, 4, 5]. Speaker representation learning captures the unique characteristics of a speaker in a compact form, and is extensively utilized across various tasks including speaker verification [3], speaker diarization [6], target speaker extraction [7], speaker-attributed automatic speech recognition [8], and other speech-related tasks [9, 10]. Researchers have studied different training schemes to learn robust speaker representations, from fully-supervised leaning [3, 1] to self-supervised learning [13, 12, 11]. With the continuous emergence of open-source projects, speech-related technology has advanced in sophistication. Table I lists some well-known open-source speech toolkits. Initially, foundational speech processing toolkits such as Kaldi [14] serve as crucial resources for both researchers and industrial applications. More recently, ESPnet [15] and SpeechBrain [16] provide novice-friendly code designs and usage. Furthermore, VoxCeleb_Trainer [17], ASV-Subtools [18], and Wespeaker [19] offer open-source platforms for the speaker community to effortlessly build models.
@@ II-B1 Visual Module
$downarrow EER()\downarrow MinDCF\downarrow$ $downarrow EER()\downarrow MinDCF\downarrow$ $downarrow EER()\downarrow MinDCF\downarrow$
@@
TABLE IV: Comparison between self-supervised learning models on VoxCeleb1-O. “*” denotes the SSL frameworks in 3D-Speaker-Toolkit.
[…]
@@ III-B Multimodal Speaker diarization
Experiments based on the proposed multimodal diarization method are conducted on a self-collected video dataset, which includes 2 to 10 speakers. Strong baselines for audio-only diarization have been established using VBx [50] and Spectral clustering(SC) [51] methods. We use E2CP integrated with spectral clustering method to leverage multimodal information to enhance speaker diarization. The comparison results are shown in Table V. Common speaker diarization metrics, including Diarization Error Rate(DER) [52], Jaccard Error Rate(JER) [53] and concatenated minimum-permutation Word Error Rate(cpWER) [54], are used. The results show that, compared to systems based solely on audio, incorporating visual or textual information both achieve notable performance improvements. Combining all three modalities of audio, visual, and textual data yields superior results on all metrics.